OpenStudy (anonymous):

Someone just answered me a 3 x equation, but im lost again, this one only has 2... a[squared]-11a+18=0

OpenStudy (anonymous):

so 2a^2-11a+18=0, correct?

OpenStudy (anonymous):

do i just not factor an a out?

OpenStudy (anonymous):

could you rewrite the equation

OpenStudy (anonymous):

the one u typed to me is correct

OpenStudy (anonymous):

is there a 2 before the first a?

OpenStudy (anonymous):

a^2-11a+18=0 or 2a^2-11a+18=0